OpenStudy (anonymous):

What is the equation I would use to figure out this question: If you have a sample of meteorite in which 10% of the Potassium -40 has decayed, how old is it?

OpenStudy (anonymous):

in this case, as the percentage of mass is given (though the half-life is needed to be given) \(\ln {\frac{m}{m_0}} = {\frac{1}{2}}^{\frac{T}{t_{\frac{1}{2}}}}\)